OpenStudy (anonymous):

Fools problem of the day: Find the last two digits of a) \(22^3+24^3+26^3+28^3 \) b) \( 23^3+24^3+25^3+26^3+27^3 \) PS:This may be too easy.

OpenStudy (zarkon):

it is ;)

OpenStudy (across):

Gotta love modular arithmetic. ;)

OpenStudy (turingtest):

I can find all the last digit easy enough, but the other one... still processing the stuff I tried to learn yesterday

OpenStudy (anonymous):

If this is just too easy: http://openstudy.com/study#/updates/4f5c5fc7e4b0602be43841dc But this ain't hard either.

OpenStudy (anonymous):

@across: I think we can solve these using just algebraic manipulations.

OpenStudy (anonymous):

Everyone just said easy, too easy and expressed their love for math but none solved it. Now someone please post the solution.

OpenStudy (anonymous):

\[ 22^3+24^3+26^3+28^3 = (22^3+28^3)+(24^3+26^3) \] \[ = 50 \times \left( ( 22^2+28^2+22\times 28) + (24^2+26^2+24\times 26) \right) \] It's easy to notice that the above is divisible by 100. Hence the last two digits are 00. Can you do the second one?